# Asymptotically tight Lagrangian dual of smooth nonconvex problems via improved error bound of Shapley-Folkman Lemma

### v1

@@ Abstract
In convex geometry, the Shapley–Folkman Lemma asserts that the nonconvexity of a Minkowski sum of $n$-dimensional bounded nonconvex sets does not accumulate once the number of summands exceeds the dimension $n$, and thus the sum becomes approximately convex. Originally published by Starr in the context of quasi-equilibrium in nonconvex market models in economics, the lemma has since found widespread use in optimization, particularly for estimating the duality gap of the Lagrangian dual of separable nonconvex problems. Given its foundational nature, we pose the following geometric question: Is it possible for the nonconvexity of the Minkowski sum of $n$-dimensional nonconvex sets to even diminish instead of just not accumulating as the number of summands increases, under some general conditions? We answer this affirmatively. First, we provide an elementary geometric proof of the Shapley–Folkman Lemma based on the facial structure of the convex hull of each set. This leads to refinement of the classical error bound derived from the lemma. Building on this new geometric perspective, we further show that when most of the sets satisfy a certain local smoothness condition which naturally arises in the epigraphs of smooth functions, their Minkowski sum converges directly to a convex set, with a vanishing nonconvexity measure. In optimization, this implies that the Lagrangian dual of block-structured smooth nonconvex problems—with potentially additional sparsity constraints—is asymptotically tight under mild assumptions, which contracts non-vanishing duality gap obtained via classical Shapley-Folkman Lemma.
@@ 1 Introduction
where $\mathcal{X}^{(i)}$ is a closed possibly nonconvex set of $\mathbb{R}^{n}$, $B^{(i)}\in\mathbb{R}^{m\times n}$ and $\mathbf{b}\in\mathbb{R}^{m}$. Although the objective and the coupling constraints in (1) appear linear, this formulation is general enough to capture nonlinear problems. Indeed, all nonlinearities can be embedded in the feasible sets $\mathcal{X}^{(i)}$ by introducing suitable auxiliary variables.
Another important program associated with (1) is the Lagrangian dual (also called Lagrangian relaxation) of (1) by dualizing the coupling constraints (1b):
$L(\lambda)$= $\inf_{\mathbf{x}}\left(\sum_{i=1}^{k}\left\langle{\mathbf{c}^{(i)}},{\mathbf{x}^{(i)}}\right\rangle\right)+\left\langle{\lambda},{B^{(i)}\mathbf{x}^{(i)}-\mathbf{b}}\right\rangle$ (2)
@@
The Lagrangian dual is computationally promising because it is always a (non-smooth) convex program and provides a valid dual bound of (1) regardless of convexity of (1): $\operatorname{\operatorname{OPT}}(\mathbf{b})\geq\operatorname{\operatorname{DUAL}}(\mathbf{b})$. On the other hand, the Lagrangian dual exploits the block-structure of (1) so that its subgradient can be obtained by solving $L(\lambda)$, which is fully decomposed into $k$ independent problems. That is, for each fixed $\lambda\geq 0$,
$B^{(i)}\mathbf{x}^{(i)}_{*}-\mathbf{b}\in\partial L(\lambda),\text{ where }\mathbf{x}_{*}:=\operatorname{\operatorname{arginf}}_{\mathbf{x}}L(\lambda).$
Therefore, given an oracle that minimizing a linear function over each $\mathcal{X}^{(i)}$, one can compute subgradient of $L(\lambda)$ and solve (3) via non-smooth optimization methods  [bagirov2014introduction, frangioni2020standard, van2018incremental]. This allows distributed algorithm to solve block-structure problems (1) and it is also known as column generation or Dantzig–Wolfe decomposition [lubbecke2010column, desaulniers2006column]. Aside from computational advantages, Lagrangian dual of nonconvex problems also finds other applications: sensitivity analysis of nonconvex problems [guo2024sensitivity], pricing in nonconvex markets [disfani2014multi, yang2019unified], market equilibrium [starr1969quasi] and online algorithm [balseiro2020dual, argue2022robust].
Many textbooks and notes state that the strong duality, that is $\operatorname{\operatorname{OPT}}(\mathbf{b})=\operatorname{\operatorname{DUAL}}(\mathbf{b})$, holds under certain regularity condition if all $\mathcal{X}^{(i)}$s are convex. This brings an impression that such level of convexity is needed for the existence of strong duality. However, the strong duality actually only requires the convexity of linear image of (1) onto the space of objective and constraints. More specifically, Let
@@
$\mathcal{P}:=\sum_{i=1}^{k}\mathcal{P}^{(i)}$
where we overload standard sum for vectors with Minkowski sum for sets and $\mathcal{P}$ is the Minkowski sum of all $\mathcal{P}^{(i)}$s. Under suitable regularity conditions, strong duality can be guaranteed whenever the image set $\mathcal{P}$ is convex [ben2001lectures]. In particular, when each set $\mathcal{X}^{(i)}$ is convex, the overall set $\mathcal{P}$ remains convex because both linear mappings and Minkowski sums preserve convexity. Therefore, the condition frequently shown in the textbooks and notes can be viewed as a sufficient condition. Although this distinction may appear subtle, it is in fact crucial—it underpins several nontrivial results concerning the strong duality in Lagrangian relaxations of nonconvex programs [polik2007survey, chandrasekaran2025lagrangian]. A notable example is the celebrated $S$-Lemma [polik2007survey], which establishes that the Lagrangian dual of a quadratic program with a single quadratic constraint is tight. This result can be interpreted as a direct consequence of Dines Lemma [dines1941mapping] that the image of two homogeneous quadratic mappings is always convex.
In this paper, we adopt this perspective and investigate the convexity of $\mathcal{P}$ to ensure the strong duality of the Lagrangian dual of (1). While each individual set $\mathcal{P}^{(i)}$ may be nonconvex, it is possible that their Minkowski sum $\mathcal{P}$ closely approximates its convex hull. This phenomenon is formally characterized by the Shapley–Folkman Lemma [starr1969quasi]. In the setting where $k\geq n$, the Shapley-Folkman Lemma states that
$\forall\mathbf{x}\in{\mathbf{conv}}(\mathcal{P}),\;\exists I\in\binom{[k]}{n}\text{ such that }\mathbf{x}\in\left(\sum_{i\in[k]\setminus I}\mathcal{P}^{(i)}\right)+\left(\sum_{i\in I}{\mathbf{conv}}(\mathcal{P}^{(i)})\right).$
Intuitively, the Shapley-Folkman Lemma implies that every point in ${\mathbf{conv}}(\mathcal{P})$ is close to some point in $\mathcal{P}$, up to at most $n$ nonconvex summands. If all $\mathcal{P}^{(i)}$ are contained within a ball of radius $R$, a quantitative measure of this approximate convexity can be obtained. In particular, under the Hausdorff distance, it has been shown [fradelizi2018convexification] that
@@
This quantitative bound can further be employed to derive an a priori estimate of the duality gap $\Delta:=\operatorname{\operatorname{OPT}}(\mathbf{b})-\operatorname{\operatorname{DUAL}}(\mathbf{b})$ associated with problem (1) (see Theorem 3).
The Shapley–Folkman lemma was originally introduced by Starr in the study of quasi-equilibria in nonconvex market models in economics. In this context, it can be interpreted as an approximate form of strong duality for Lagrangian relaxations. Its connection to optimization was later developed in works such as [bertsekas1983optimal, aubin1976estimates, udell2016bounding], which used the lemma to quantify the duality gap of problem (1). More recently, several works have focused on refining the quantitative bounds provided by the Shapley–Folkman lemma, particularly on improving estimates of the Hausdorff distance $d_{H}(\mathcal{P},{\mathbf{conv}}(\mathcal{P}))$ or similar notion of nonconvexity; see, for example, [bi2016refined, kerdreux2023stable, nguyen2022near]. An analogue of the Shapley–Folkman lemma for discrete convex sets has also been developed in [murota2025shapley]
In this work, we seek to refine the quantitative characterization of the Hausdorff distance $d_{H}(\mathcal{P},{\mathbf{conv}}(\mathcal{P}))$, given its central role in assessing the tightness of Lagrangian duals. We begin by presenting a new and elementary geometric proof of the Shapley-Folkman Lemma, which leverages the facial structure of each $\operatorname{conv}(\mathcal{P}^{(i)})$. This perspective enables us to refine the classical error bound. Building on this geometric insight, we further demonstrate that the Minkowski sum of the sets $\mathcal{P}^{(i)}$ can directly converges to its convex hull as the number of summands grows, i.e., $d_{H}(\mathcal{P},\operatorname{conv}(\mathcal{P}))\to 0$ as $k\to\infty$, provided enough sets satisfy some mild “local smoothness” condition (see Theorem 4.19). Such smoothness naturally arises in the epigraphs of smooth (possibly nonconvex) functions, allowing us to establish that the following problem admits an asymptotically tight Lagrangian dual as $k\to\infty$:
@@
where each $f^{(i)}$ is a smooth (possibly nonconvex) function, and $B^{(i)}$ and $s^{(i)}$ satisfy some mild conditions (see Theorem 5.31). Note that we only requires that the sparsity level $s^{(i)}$ is not too small, i.e., $s^{(i)}\geq m+1,\forall i\in[k]$. To model problems without sparsity constraints, one may simply let $s^{(i)}=n,\forall i\in[k]$. To the best of our knowledge, this is the first result establishing an asymptotically vanishing nonconvexity measures for Minkowski sums of nonconvex sets within the framework of the Shapley–Folkman lemma.
This work focuses on Minkowski sum of $\mathcal{P}^{(i)}$ without averaging since its convexity directly reflects the duality gap of problem (1). In the classical applications of the Shapley–Folkman Lemma, the analysis is typically performed on the Minkowski average of the sets $\mathcal{P}^{(i)}$. This choice arises because the the error bound $d_{H}(\mathcal{P},{\mathbf{conv}}(\mathcal{P}))$ does not vanish in the classic analysis of Shapley-Folkman Lemma but $d_{H}(\mathcal{P},{\mathbf{conv}}(\mathcal{P}))$ is both homogeneous and independent of $k$, which immediately yields
@@
2 Preliminary
In this section, we introduce several useful notation and results to study Shapley-Folkman Lemma and the Lagrangian relaxation of (1).
@@ 2.1 Notation
For a set $A\subset\mathbb{R}^{d}$, we define its radius as $\operatorname{\operatorname{R}}(A):=\inf_{x\in\mathbb{R}^{d}}\left\{r\geq 0:A\subseteq\mathcal{B}(x,r)\right\}$, where $\mathcal{B}(x,r):=\{y\in\mathbb{R}^{d}:\|y-x\|\leq r\}$ denotes the closed Euclidean ball of radius $r$ centered at $x$. Given two sets $A,B\subset\mathbb{R}^{d}$, their Minkowski sum is defined as $A+B:=\{a+b:a\in A,\;b\in B\}.$ For any scalar $\lambda\geq 0$, the Minkowski scaling of $A$ is $\lambda A:=\{\lambda a:a\in A\}.$We denote by ${\mathbf{conv}}(A):=\left\{\sum_{i=1}^{m}\lambda_{i}a_{i}\;\middle|\;a_{i}\in A,\ \lambda_{i}\geq 0,\ \sum_{i=1}^{m}\lambda_{i}=1\right\}$ the convex hull of $A$. The closed convex hull is defined as $\overline{{\mathbf{conv}}}(A):=\operatorname{cl}({\mathbf{conv}}(A)),$ where $\operatorname{cl}(A)$ denotes the topological closure of $A$.
@@ 2.2 Nonconvexity measure and convex hull operator
Given a set $A\subseteq\mathbb{R}^{n}$, the exists a rich body of literature on how to measure non-convexity of $A$ due to its wide-ranging applications in nonconvex optimization. Throughout this paper, we will focus on two measures:
@@
    (Hausdorff distance from the convex hull) Given a closed convex set $K$ with 0 in its interior, we define
@@
    (The inner radius of a nonconvex set [starr1969quasi])
    $\Xi(A):=\sup_{\mathbf{x}\in{\mathbf{conv}}(A)}\{\inf{\operatorname{\operatorname{R}}(T):T\subseteq A,\mathbf{x}\in{\mathbf{conv}}(T})\}$
@@
In the literature on the Shapley–Folkman Lemma, several measures of nonconvexity have been proposed. The measure $\Xi(\cdot)$, introduced in the seminal work by Starr [starr1969quasi], is used to establish the notion of quasi-equilibrium in nonconvex market models in economics. In this paper, we focus primarily on $\Phi(\cdot)$, as it is more directly related to the duality gap in Lagrangian dual (see Theorem 3). Other nonconvexity measures, exhibiting distinct behaviors and intended for different analytical purposes, are not listed in this paper; for a comprehensive overview, we refer readers to the survey [fradelizi2018convexification].
@@
[wegmann1980einige, fradelizi2018convexification] Let $A$ be a subset of $\mathbb{R}^{n}$, then
@@
Claim 1.
[fradelizi2018convexification] Given a closed bounded convex set $K$, $\Phi^{K}(\cdot)$ are subadditive and positive homogeneous. That is for any $A,B\subseteq\mathbb{R}^{n}$ and $\kappa>0$, it follows that
$\Phi^{K}(A+B)\leq\Phi^{K}(A)+\Phi^{K}(B);$
$\Phi^{K}(\kappa A)=\kappa\Phi^{K}(A).$
Similar results hold for $\Xi^{2}(\cdot)$.
@@
Claim 2.
($\mathbf{conv}$} commutes with Minkowski sum) Let $A_{1},\dots,A_{k}$ be nonempty sets of $\mathbb{R}^{n}$. Then it follows that
@@
Claim 3.
($\mathbf{conv}$} commutes with linear projection)
${\mathbf{conv}}(\mathcal{P}^{(i)})=\left\{(t,\mathbf{d})\;\middle|\;\begin{array}[]{@{}l@{}}\exists\mathbf{x}\in{\mathbf{conv}}\left(\mathcal{X}^{(i)}\right),t=\left\langle{\mathbf{c}^{(i)}},{\mathbf{x}^{(i)}}\right\rangle,\mathbf{d}=B^{(i)}\mathbf{x}^{(i)}\end{array}\right\}.$
3 Shapley-Folkman Lemma and duality gap of Lagrangian relaxation
In this section, we state Shapley-Folkman Lemma formally and show how it is related to duality gap of Lagrangian relaxation.

### v2

@@ Abstract
The Shapley–Folkman Lemma, a foundational result in convex geometry, asserts that the nonconvexity of a Minkowski sum of $n$-dimensional bounded nonconvex sets does not accumulate once the number of summands exceeds the dimension $n$, and thus the sum becomes approximately convex. This Lemma, originally published by Starr is used to show quasi-equilibrium in nonconvex market models in economics and equivalently used to estimate the duality gap of the Lagrangian dual of separable nonconvex problems. Given its foundational nature, we pose the following geometric question: Is it possible for the nonconvexity of the Minkowski sum of $n$-dimensional nonconvex sets to vanish asymptotically, rather than merely converge to a constant, as the number of summands increases? We answer this question affirmatively. First, we provide an elementary geometric proof of the Shapley–Folkman Lemma based on the facial structure of the convex hull of each set under a pointedness assumption. This leads to a small refinement on the classical error bound derived from the Lemma in the pointed case. Building on this new geometric perspective, we next show that if the sets satisfy a certain local smoothness condition, which naturally arises, for example, in the epigraphs of smooth functions, then their Minkowski sums converge directly to a convex set as the nonconvexity measure vanishes asymptotically. This result has direct implications for optimization and economic theory. In particular, we show that the Lagrangian dual of block-structured smooth nonconvex problems, possibly subject to additional sparsity constraints, is asymptotically tight under mild assumptions. This sharply improves upon classical Shapley–Folkman based bounds, which generally predict a nonvanishing duality gap. Equivalently, from an economic perspective, this result implies that, asymptotically as the size of the economy increases, competitive equilibrium exists with smooth utility functions, notwithstanding underlying nonconvexities.
keywords:
Shapley–Folkman Lemma, Strong Lagrangian Dual, Competitive Equilibrium
pacs:
Mathematics Subject Classification]90C11 $\cdot$ 90C46
@@ 1 Introduction
where $\mathcal{X}^{(i)}$ is a closed possibly nonconvex set of $\mathbb{R}^{n}$, $B^{(i)}\in\mathbb{R}^{m\times n}$ and $\mathbf{b}\in\mathbb{R}^{m}$. Although the objective and the coupling constraints in (1) are linear, this formulation is general enough to capture many nonlinearities, by embedding these into the feasible sets $\mathcal{X}^{(i)}$, potentially through the introduction of suitable auxiliary variables.
An important relaxation of (1) is the Lagrangian dual of (1), obtained by dualizing the coupling constraints (1b) across various blocks:
$L(\lambda)$= $\inf_{\mathbf{x}}\left(\sum_{i=1}^{k}\left\langle{\mathbf{c}^{(i)}},{\mathbf{x}^{(i)}}\right\rangle+\left\langle{\lambda},{B^{(i)}\mathbf{x}^{(i)}}\right\rangle\right)-\left\langle{\lambda},{\mathbf{b}}\right\rangle$ (2)
@@
The Lagrangian dual of the block-structured problem (1) plays a central role in both optimization and economic theory. From an optimization perspective, dualizing the coupling constraints yields a convex (possibly nonsmooth) program (3) regardless of the nonconvexity of the feasible sets $\mathcal{X}^{(i)}$, thereby providing a tractable and globally valid lower bound on $\operatorname{\operatorname{OPT}}(\mathbf{b})$. Moreover, the dual formulation exploits the separable block structure of the primal problem: for any fixed multiplier $\lambda\geq 0$, the Lagrangian $L(\lambda)$ decomposes into independent subproblems across blocks, enabling scalable algorithms based on decomposition, subgradient methods, and price-based coordination [1, 2, 3]. From an economic viewpoint, the dual variables $\lambda$ admit a natural interpretation as equilibrium prices associated with the shared resource constraint $\sum_{i=1}^{k}B^{(i)}x^{(i)}\leq\mathbf{b}$ [4]. In this interpretation, each block represents an individual agent or firm that optimizes its own objective given prices $\lambda$, while the dual objective aggregates these decentralized decisions and enforces market clearing. Consequently, the duality gap between $\operatorname{\operatorname{OPT}}(\mathbf{b})$ and $\operatorname{\operatorname{DUAL}}(\mathbf{b})$ quantifies the extent to which a competitive equilibrium exists or fails to exist in the presence of nonconvexities, making the Lagrangian framework a unifying bridge between large-scale optimization and equilibrium analysis in economic theory.
One of the most desirable properties one aims to establish for the Lagrangian dual, in both optimization and economic theory, is strong duality. In optimization, strong duality certifies that the optimal value of a (possibly relaxed) dual problem coincides with that of the primal, thereby justifying convex relaxations, decomposition methods, and price-based algorithms. In economic theory, strong duality corresponds to the existence of supporting prices that decentralize a socially optimal allocation, that is, establishing the existence of a competitive equilibrium.
Many textbooks state that strong duality, that is $\operatorname{\operatorname{OPT}}(\mathbf{b})=\operatorname{\operatorname{DUAL}}(\mathbf{b})$, holds under certain regularity conditions when all $\mathcal{X}^{(i)}$s are convex. However, strong duality actually only requires the convexity of linear image of (1) onto the space of objective and constraints [5]. More specifically, let
@@
$\mathcal{P}:=\sum_{i=1}^{k}\mathcal{P}^{(i)},$
where we overload standard sum for vectors with Minkowski sum for sets and $\mathcal{P}$ is the Minkowski sum of all $\mathcal{P}^{(i)}$s. Under suitable regularity conditions, strong duality can be guaranteed whenever the image set $\mathcal{P}$ is convex [5]. In particular, when each set $\mathcal{X}^{(i)}$ is convex, the overall set $\mathcal{P}$ remains convex because both linear mappings and Minkowski sums preserve convexity. Therefore, the condition frequently shown in textbooks can be viewed as a sufficient condition. Although this distinction may appear subtle, it is in fact crucial—it underpins several nontrivial results concerning the strong duality in Lagrangian relaxations of nonconvex programs [6, 7]. A notable example is the celebrated $S$-Lemma [6], which establishes that the Lagrangian dual of a quadratic program with a single quadratic constraint is tight. This result can be interpreted as a direct consequence of Dines Lemma [8] that the image of two homogeneous quadratic mappings is always convex.
In this paper, we adopt this perspective and investigate the convexity of $\mathcal{P}$ to ensure the strong duality of the Lagrangian dual of (1). While each individual set $\mathcal{P}^{(i)}$ may be nonconvex, it is possible that their Minkowski sum $\mathcal{P}$ closely approximates its convex hull. This phenomenon is formally characterized by the Shapley–Folkman Lemma [4]. In the setting where $k\geq n$, the Shapley-Folkman Lemma states that
$\forall\mathbf{x}\in{\mathbf{conv}}(\mathcal{P}),\;\exists I\subseteq[k],|I|\leq n\text{ such that }\mathbf{x}\in\left(\sum_{i\in[k]\setminus I}\mathcal{P}^{(i)}\right)+\left(\sum_{i\in I}{\mathbf{conv}}\left(\mathcal{P}^{(i)}\right)\right).$
Here $k]=\{1,\dots,k\}$. Intuitively, the Shapley-Folkman Lemma implies that every point in ${\mathbf{conv}}(\mathcal{P})$ is close to some point in $\mathcal{P}$, up to at most $n$ ‘nonconvex summands’. If all $\mathcal{P}^{(i)}$s are contained within a ball of radius $R$, then a quantitative measure of this approximate convexity can be obtained. In particular, under the Hausdorff distance with $l_{2}$ norm, it has been shown [9] that
@@
This quantitative bound can be employed to derive an a priori estimate of the duality gap $\Delta:=\operatorname{\operatorname{OPT}}(\mathbf{b})-\operatorname{\operatorname{DUAL}}(\mathbf{b})$ associated with problem (1) (see Theorem 3 in Section 3).
The Shapley–Folkman lemma was originally introduced by Starr [4] in the study of quasi-equilibria in nonconvex market models in economics. In this context, it can be interpreted as an approximate form of strong duality for Lagrangian duals. Its connection to optimization was later developed in works such as [10, 11, 12], which used the lemma to quantify the duality gap of problem (1). More recently, several works have focused on refining the quantitative bounds provided by the Shapley–Folkman lemma, particularly on improving estimates of the Hausdorff distance $d_{H}(\mathcal{P},{\mathbf{conv}}(\mathcal{P}))$ or similar notion of nonconvexity; see, for example, [13, 14, 15]. An analog of the Shapley–Folkman lemma for discrete convex sets has also been developed in [16].
In this work, we seek to refine the quantitative characterization of the Hausdorff distance $d_{H}(\mathcal{P},{\mathbf{conv}}(\mathcal{P}))$, given its central role in assessing the tightness of Lagrangian duals. We begin by presenting a new and elementary geometric proof of the Shapley-Folkman Lemma, which leverages the facial structure of each $\operatorname{conv}(\mathcal{P}^{(i)})$. This perspective enables us to refine the classical error bound under a pointedness assumption (see Theorem 4.16 in Section 4). Building on this geometric insight, we further demonstrate that the Minkowski sum of the sets $\mathcal{P}^{(i)}$ can directly converge to its convex hull as the number of summands grows, i.e., $d_{H}(\mathcal{P},\operatorname{conv}(\mathcal{P}))\to 0$ as $k\to\infty$, provided enough sets satisfy some mild “local smoothness” condition (see Theorem 4.22 in Section 4). To the best of our knowledge, this is the first result establishing an asymptotically vanishing nonconvexity measure for Minkowski sums of nonconvex sets within the framework of the Shapley–Folkman Lemma.
Such smoothness naturally arises in the epigraphs of smooth (possibly nonconvex) functions, allowing us to establish that the following problem admits an asymptotically tight Lagrangian dual as $k\to\infty$:
@@
where each $f^{(i)}$ is a smooth (possibly nonconvex) function, $B^{(i)}$s satisfy some mild conditions, and $s^{(i)}$ is not too small, i.e., $s^{(i)}\geq m,\forall i\in[k]$ where $B^{(i)}$ has $m$ rows (see Theorem 5.38 in Section 5) . To model problems without sparsity constraints, one may simply drop these constraints, that is let $s^{(i)}=\textup{number of variables in }\mathbf{x}^{i}$, for all $i\in[k]$.
Finally, we note here that this work focuses on Minkowski sum of $\mathcal{P}^{(i)}$s without averaging since its convexity directly reflects the duality gap of problem (1). In the classical applications of the Shapley–Folkman Lemma, the analysis is typically performed on the Minkowski average of the sets $\mathcal{P}^{(i)}$. This choice arises because the error bound $d_{H}(\mathcal{P},{\mathbf{conv}}(\mathcal{P}))$ does not vanish in the classic analysis of Shapley-Folkman Lemma but $d_{H}(\mathcal{P},{\mathbf{conv}}(\mathcal{P}))$ is both homogeneous and independent of $k$, which immediately yields
@@
The remainder of the paper is organized as follows. In Section 2, we present standard notation and preliminary results. Section 3 establishes a connection between $d_{H}(\mathcal{P},{\mathbf{conv}}(\mathcal{P}))$ and $\Delta:=\operatorname{\operatorname{OPT}}(\mathbf{b})-\operatorname{\operatorname{DUAL}}(\mathbf{b})$ in Theorem 3. Section 4 presents a new proof of Shapley-Folkman Lemma under the assumption of pointedness of the convex hull of the non-convex sets (see Lemma 4.7 and discussion following it) which allows us to establish a slightly better bound on $d_{H}(\mathcal{P},{\mathbf{conv}}(\mathcal{P}))$, in Theorem 4.16. Our main result of showing that $d_{H}(\mathcal{P},{\mathbf{conv}}(\mathcal{P}))$ goes to zero as $k$ goes to $\infty$ under the presence of a local smoothness property, is presented in this section in Theorem 4.22. Finally, in Section 5, we establish the strong duality result stated informally above, formalized as Theorem 5.38.
2 Preliminary Results and Notation
In this section, we introduce several useful notation and results to study the Shapley-Folkman Lemma and the Lagrangian dual of (1).
@@ 2.1 Notation
For a set $A\subset\mathbb{R}^{d}$, we define its radius as $\operatorname{\operatorname{R}}(A):=\inf_{x\in\mathbb{R}^{d}}\left\{r\geq 0:A\subseteq\mathcal{B}(x,r)\right\}$, where $\mathcal{B}(x,r):=\{y\in\mathbb{R}^{d}:\|y-x\|_{2}\leq r\}$ denotes the closed Euclidean ball of radius $r$ centered at $x$. Given two sets $A,B\subset\mathbb{R}^{d}$, their Minkowski sum is defined as $A+B:=\{a+b:a\in A,\;b\in B\}.$ For any scalar $\lambda\geq 0$, the Minkowski scaling of $A$ is $\lambda A:=\{\lambda a:a\in A\}.$We denote by ${\mathbf{conv}}(A):=\left\{\sum_{i=1}^{m}\lambda_{i}a_{i}\;\middle|\;a_{i}\in A,\ \lambda_{i}\geq 0,\ \sum_{i=1}^{m}\lambda_{i}=1\right\}$ the convex hull of $A$. The closed convex hull is defined as $\overline{{\mathbf{conv}}}(A):=\operatorname{cl}({\mathbf{conv}}(A)),$ where $\operatorname{cl}(A)$ denotes the topological closure of $A$. The relative boundary and the recession cone of a closed convex set $C$ are denoted using $\textup{relbd}(C)$ and $\text{rec}(C)$. The kernel of a linear map $\mathcal{A}$ is denoted by $\ker(\mathcal{A})$. We use $\operatorname{\operatorname{epi}}(f)$ to denote the epigraph of a function $f$.
@@ 2.2 Nonconvexity measure and convex hull operator
For a given set $A\subseteq\mathbb{R}^{n}$, there exists a rich body of literature on how to measure non-convexity of $A$ due to its wide-ranging applications in nonconvex optimization. Throughout this paper, we will focus on two measures:
@@
    Hausdorff distance from the convex hull: Given a closed convex set $K$ with 0 in its interior, we define
@@
    The inner radius of a nonconvex set [4]:
    $\Xi(A):=\sup_{\mathbf{x}\in{\mathbf{conv}}(A)}\left\{\inf{\operatorname{\operatorname{R}}(T):T\subseteq A,\mathbf{x}\in{\mathbf{conv}}(T})\right\}.$
@@
The measure $\Xi(\cdot)$, introduced in the seminal work by Starr [4], is used to establish the notion of quasi-equilibrium in nonconvex market models in economics. In this paper, we focus primarily on $\Phi(\cdot)$, as it is more directly related to the duality gap in Lagrangian dual (see Theorem 3). See Figure 2 and 2 for illustration of the $\Phi(\cdot)$ and $\Xi(\cdot)$ nonconvexity measure. Other nonconvexity measures, exhibiting distinct behaviors and intended for different analytical purposes, are not listed in this paper; for a comprehensive overview, we refer readers to the survey [9]. We present next a standard result on these non-convexity measures.
@@
[17, 9] Let $A$ be a subset of $\mathbb{R}^{n}$, then
@@
Lemma 2.
[9] ($\mathbf{conv}$} commutes with Minkowski sum) Let $A_{1},\dots,A_{k}$ be nonempty sets of $\mathbb{R}^{n}$. Then it follows that
@@
Lemma 3.
[18] ($\mathbf{conv}$} commutes with linear projection)
${\mathbf{conv}}\left(\mathcal{P}^{(i)}\right)=\left\{(t,\mathbf{d})\;\middle|\;\begin{array}[]{@{}l@{}}\exists\mathbf{x}\in{\mathbf{conv}}\left(\mathcal{X}^{(i)}\right),t=\left\langle{\mathbf{c}^{(i)}},{\mathbf{x}^{(i)}}\right\rangle,\mathbf{d}=B^{(i)}\mathbf{x}^{(i)}\end{array}\right\}.$